OpenStudy (anonymous):

A public library has fiction and nonfiction books in the ratio 3:2 . If the library 360 has fiction books, how many nonfiction books does it have?

OpenStudy (anonymous):

try to multiply ratios

OpenStudy (anonymous):

fiction/nonfiction=3:2 360/nonfiction=3/2 360*2/3=nonfiction nonfiction=240

OpenStudy (anonymous):

nader1 is right

OpenStudy (anonymous):

Fict = 3 Nonfict = 2 thus, 360/3 = 120 ( 1 unit is 120) 120*2 = 240

OpenStudy (anonymous):

this for more detail :) for you to understand :)

OpenStudy (anonymous):

understand jakie?